# Deep Reinforcement Learning for Robotic Bipedal Locomotion: A Brief Survey

### v5

@@ Abstract
Bipedal robots are gaining global recognition due to their potential applications and advancements in artificial intelligence, particularly through Deep Reinforcement Learning (DRL). While DRL has significantly advanced bipedal locomotion, the development of a unified framework capable of handling a wide range of tasks remains an ongoing challenge. This survey systematically categorises, compares, and analyses existing DRL frameworks for bipedal locomotion, organising them into end-to-end and hierarchical control schemes. End-to-end frameworks are evaluated based on their learning approaches, while hierarchical frameworks are examined in terms of layered structures that integrate learning-based or traditional model-based methods. We provide a detailed evaluation of the composition, strengths, limitations, and capabilities of each framework. Additionally, this survey identifies key research gaps and proposes future directions aimed at creating a more integrated and efficient unified framework for bipedal locomotion, with wide-ranging applications in real-world environments.
@@ I Introduction
Humans navigate complex environments and perform diverse locomotion tasks with remarkable efficiency using just two legs. Bipedal robots, which closely mimic human form, have a distinct advantage over wheeled or tracked alternatives, particularly when traversing uneven and challenging terrains. Furthermore, bipedal humanoid robots are specifically designed to function in human-centric environments, enabling seamless interaction with tools and infrastructure intended for human use. This makes them highly adaptable to a wide range of tasks in such settings.
@@
As a result, bipedal robots hold significant potential for real-world applications [1]. In manufacturing, they can perform tasks efficiently without requiring additional tools, thereby enhancing productivity and reducing labour demands [2, 3, 4]. Their agility is particularly advantageous in complex environments, such as multi-level workplaces. Bipedal robots are also well-suited for tasks that involve the use of human-designed tools, making them valuable for assisting in daily tasks, healthcare, and rehabilitation [5]. Furthermore, they show considerable promise in search and rescue operations, where they can navigate hazardous and unpredictable terrains [6, 7, 8].
Figure 1: Common bipedal and humanoid robots used as platforms for testing DRL frameworks. (a) NAO, a toy-like 3D humanoid robot, actuated by servo motors [9]. (b) Rabbit, a 2D bipedal robot, actuated by torque control [10]. (c) Cassie, a 3D bipedal robot, also actuated by torque control [11]. (d) ATLAS, a 3D humanoid robot, driven by hydraulics [12]. (e) Digit, a full human-sized 3D humanoid robot, an upgrade based on Cassie and actuated by torque control [13].
@@
Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [14, 15, 16]. Early methods, such as the Linear Inverted Pendulum Model (LIPM) [17], allowed for simplified representations of the dynamics involved in bipedal movement, enabling easier analysis and control. As research progressed, full dynamic models were introduced to better capture the complexities of real-world locomotion. Advanced methods such as Model Predictive Control (MPC) [18, 19] and Trajectory Optimisation (TO) [20, 21, 14] exploit a predefined dynamics model to solve constrained optimal-control problems planning footsteps, CoM motion, and contact forces. While model-based approaches offer rapid convergence and predictive capabilities, they often fall short in dynamically complex and uncertain environments, where adaptability is key.
@@
Reinforcement Learning (RL)-based methods, particularly Deep Reinforcement Learning (DRL), are effective in optimising robot control policies through direct interaction with the environment [22], which offers a distinct advantage. Unlike model-based approaches, which rely on predefined dynamics and can struggle with unforeseen scenarios, DRL enables robots to autonomously discover control strategies through trial and error, allowing for greater adaptability and robustness in diverse environments. Additionally, hybrid methods that combine model-based and learning-based techniques further enhance planning and control by leveraging the strengths of both approaches.
Despite these advancements, a key question remains: Is there a unified framework capable of enabling bipedal robots to effectively manage a diverse range of locomotion tasks? If not, how should such a framework be designed, and what features or components should it include?
To provide a clear benchmark for answering this question, we first define a unified framework as a control architecture capable of supporting consistent learning, transfer, and execution across arbitrary robot morphologies, terrains, and task definitions. The ultimate aim is to develop genuinely generalist locomotion policies that achieve robust zero-shot generalisation to previously unseen scenarios, embodiments, and objectives.
With this definition as our guide, this survey examines recent advancements in DRL-based frameworks, categorising control schemes into two primary types: (i) end-to-end and (ii) hierarchical. These frameworks enable robots to generate diverse gaits, including walking, running, and turning, while also allowing them to navigate complex terrains, such as stairs, slopes, and deformable surfaces.
@@
End-to-end frameworks directly map robot states to joint-level control outputs, while hierarchical frameworks decompose decision-making into multiple layers. In hierarchical systems, a High-Level (HL) planner governs navigation and path planning, while a Low-Level (LL) controller handles fundamental locomotion tasks. The task-level decision-making tier interfaces directly with user commands or predefined tasks, forming a structured approach to robotic control.
The evolution of RL in bipedal robotics has spurred a dynamic growth in innovative applications. Although the application of RL to simple 2D bipedal robots began in 2004 [23, 24], it took several years before DRL algorithms emerged. These DRL-based methods have since shown promising results in physical simulators [25, 26, 27]. Agility Robotics introduced the first sim-to-real end-to-end learning framework in 2019, which was applied on the 3D torque-controlled bipedal robot Cassie, as shown in Fig. 1(c) [11]. Besides model-based reference learning, the policy can also incorporate motion capture data [28, 29, 30, 31], or start from scratch [32] to explore solutions freely. Recent studies demonstrate that end-to-end frameworks robustly handle complex and diverse tasks [33, 34, 35].
Similarly, hierarchical structures have garnered significant interest. Within this subset, the hybrid approach combines RL-based and model-based methods to enhance both planning and control strategies. A notable framework employs a learned HL planner coupled with a LL model-based controller, often referred to as the Cascade-structure or Deep Planning Hybrid Scheme [21, 36, 37]. Another innovative construction integrates a learned feedback controller with an HL planner, categorised under the DRL Feedback Control Hybrid Scheme [38, 39]. Additionally, a learned hierarchical control scheme [40] decomposes locomotion into various tasks, focusing each layer on specific functions such as navigation and fundamental locomotion skills [25, 26, 41]. In order to provide a clearer understanding of the current landscape, we have categorised existing DRL frameworks, as depicted in Fig. 2.
@@
Figure 2: Classification of DRL-based control schemes. The approaches are broadly categorised into two main paradigms: End-to-end frameworks, which learn a single policy from sensory inputs to motor commands, and Hierarchical frameworks, which decompose the control problem into multiple levels. Within the end-to-end paradigm, a key distinction is between Reference-free learning (learning from scratch) and Reference-based learning (tracking a predefined motion). Hierarchical structures include Hybrid control schemes, which synergistically combine learned components with traditional model-based controllers.
Current progress shows that we are still far from achieving a unified framework, despite two mainstream types of framework structure. Although bipedal locomotion tasks have evolved significantly in the past decade. Earlier frameworks primarily focused on basic walking stabilisation and simple terrain adaptation, whereas recent developments have introduced complex tasks such as dynamic stair climbing, parkour-inspired locomotion, and loco-manipulation. Notably, the DARPA Robotics Challenge (DRC) has established a structured benchmark for bipedal humanoid locomotion and manipulation [42], emphasizing real-world deployment and robustness.
@@
Although several reviews discuss RL for general robotics [22] and model-based methods for bipedal robots [14, 15, 16], none specifically focus on DRL-based frameworks for bipeds. To address this gap, this survey reviews relevant literature by applying the following selection criteria: (1) studies that examine DRL frameworks specifically for bipedal robots, (2) research involving both simulated and physical bipedal robots, (3) approaches that improve policy transfer from simulation to real-world environments, and (4) publications from the last five years (2018 - April 2024) sourced from reputable databases and conferences, including Google Scholar, IEEE Xplore, Web of Science, ArXiv, and major robotics conferences such as CoRL, RSS, ICRA, IROS, and Humanoids.
@@
The search was conducted using the key terms “Deep Reinforcement Learning” or “Reinforcement Learning” in combination with “Bipedal Locomotion”, “Bipedal Walking”, “Biped Robot”, “Humanoid Robot”, or “Legged Robot”. The most relevant and impactful works were manually selected for further review. This survey is intended for readers with a foundational background in robotics who are transitioning to DRL methods, providing an overview of a wide range of approaches with simplified explanations where necessary. For clarity, in this survey, the term “humanoid robot” refers specifically to anthropomorphic bipedal robots.
@@
  • A detailed comparison of each control scheme, highlighting their strengths, limitations, and characteristics.
@@
The survey is organised as follows: Section II focuses on end-to-end frameworks, categorised by learning approaches. Section III details hierarchical frameworks, classified into three main types. Section IV presents specific limitations and challenges, highlighting their connection to the previously surveyed content. Section V addresses potential future pathways, opportunities, and two proposed conceptual models that extend the end-to-end and hierarchical paradigms. Finally, Section VI concludes the survey.
@@ II End-to-end framework
The end-to-end DRL framework represents a holistic approach where a single neural network (NN) policy, denoted $\pi(\cdot):\mathcal{X}\rightarrow\mathcal{U}$, directly maps sensory inputs $\mathcal{X}$, such as images, lidar data, or proprioceptive feedback [43], along with user commands [32] or pre-defined references [44], into joint-level control actions $\mathcal{U}$. Here, $\mathcal{X}$ represents the sensory input space, $\mathcal{U}$ refers to the space of control actions, and $\pi(\cdot)$ denotes the policy function. This framework obviates the need for manually decomposing the problem into sub-tasks, streamlining the control process.
End-to-end strategies primarily simplify the design of LL tracking to basic elements, such as a Proportional-Derivative (PD) controller. These methods can be broadly categorised based on their reliance on prior knowledge into two types: reference-based and reference-free. The locomotion skills developed through these diverse learning approaches exhibit considerable variation in performance and adaptability.
In the following sections, we will delve into various representation frameworks, exploring their characteristics, limitations, and strengths in comprehensive detail. To facilitate an understanding of these distinctions, Table I provides a succinct overview of the frameworks discussed.
TABLE I: Summary and comparison of reference-based and reference-free Learning approaches for end-to-end framework. The dashed line in the implementation flow chat refers to optional.00footnotetext: Forward Walk involves the bipeds walking straight ahead. Unidirectional Walk enables the bipeds to move both forward and backward within a range of desired velocities. Omni-Walk grants the bipeds the ability to walk in any direction. Versatile Walk allows the bipeds to walk forward, backward, turn, and move sideways, providing extensive movement capabilities. Periodic Motions entails the execution of various repeated gait patterns, such as walking, hopping, or galloping. Versatile Jump refers to jump towards different desired targets. Versatile Motions cover performing a broad array of motions, both periodic and aperiodic such as jumping.

### v6

@@ Abstract
Bipedal robots are gaining global recognition due to their potential applications and the rapid advancements in artificial intelligence, particularly through Deep Reinforcement Learning (DRL). While DRL has significantly advanced bipedal locomotion, the development of a unified framework capable of handling a wide range of tasks remains an ongoing challenge. This survey systematically categorises, compares, and analyses existing DRL frameworks for bipedal locomotion, organising them into end-to-end and hierarchical control schemes. End-to-end frameworks are evaluated based on their learning approaches, whereas hierarchical frameworks are examined in terms of their layered structures that integrate learning-based and traditional model-based methods. We provide a detailed evaluation of the composition, strengths, limitations, and capabilities of each framework. Furthermore, this survey identifies key research gaps and proposes future directions aimed at creating a more integrated and efficient unified framework for bipedal locomotion, with broad applicability in real-world environments.
@@ I Introduction
Humans navigate complex environments and perform diverse locomotion tasks with remarkable efficiency using only two legs. Bipedal robots, which closely mimic the human form, possess distinct advantages over wheeled or tracked alternatives, particularly when traversing uneven and challenging terrains. Furthermore, bipedal humanoid robots are specifically designed to operate in human-centric environments, enabling seamless interaction with tools and infrastructure intended for human use. This makes them highly adaptable to a wide range of tasks in such settings.
Figure 1: Representative bipedal and humanoid robots illustrating the diversity of platforms for locomotion research and development. (a) Cassie: a torque-controlled bipedal robot designed for agile locomotion. (b) Digit: a full-sized humanoid robot evolved from Cassie and actuated by torque control. (c) H1: a full-size, electric, torque-controlled humanoid robot developed by Unitree Robotics. (d) G1: a compact humanoid robot from Unitree featuring lightweight design and high joint backdrivability. (e) Atlas: a fully electric humanoid robot developed by Boston Dynamics.
@@
As a result, bipedal robots hold significant potential for real-world applications [2024_tong_advancements_review]. In manufacturing, they can perform tasks efficiently without requiring additional tools, thereby enhancing productivity and reducing labour demands [industrial_robotics_applications, collaborative_industrial, 2023_vidoeo_demo_digit_locomaniputlation]. Their agility is particularly advantageous in complex environments such as multi-level workplaces. Bipedal robots are also well suited to tasks that involve the use of human-designed tools, making them valuable for assisting in daily activities, healthcare, and rehabilitation [2019_bingjing_human_exoskeleton]. Moreover, they show considerable promise in search-and-rescue operations, where they can navigate hazardous and unpredictable terrains [2015_underwaterrobot_review, quadSpace1, quadSpace2].
@@
Traditional approaches to bipedal locomotion control, such as model-based methods, have been prevalent since the 1980s [2017_biped_model-based_method_review, 2021_Jenna_bipedrobot_review, 2021_Justin_review_recentporgressinleggedrobot]. Early methods, such as the Linear Inverted Pendulum Model (LIPM) [2013_3DSLIPmodel_running], provided simplified representations of the dynamics involved in bipedal motion, enabling easier analysis and control. As research progressed, full dynamic models were introduced to better capture the complexities of real-world locomotion. Advanced methods such as Model Predictive Control (MPC) [2022_Hou_drlasparametertoMPC, 2023_junheng_steppingstones_MPCWBC] and Trajectory Optimisation (TO) [2016_LIPM_trajectoryoptimization, 2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy, 2017_biped_model-based_method_review] exploit predefined dynamic models to solve constrained optimal-control problems that plan footsteps, centre-of-mass (CoM) motion, and contact forces. While model-based approaches offer rapid convergence and predictive capabilities, they often struggle in dynamically complex and uncertain environments where adaptability is essential.
@@
Reinforcement learning (RL)-based methods, particularly deep reinforcement learning (DRL), are effective in optimising robot control policies through direct interaction with the environment [2020_systematic_review], which provides a distinct advantage. Unlike model-based approaches, which rely on predefined dynamics and may fail under unforeseen conditions, DRL enables robots to autonomously discover control strategies through trial and error, achieving greater adaptability and robustness in diverse environments. In addition, hybrid methods that combine model-based and learning-based techniques further enhance planning and control by leveraging the strengths of both paradigms.
Despite these advancements, research in DRL-based locomotion remains highly fragmented, with inconsistencies in training pipelines, reward formulations, observation spaces, and evaluation setups that hinder systematic benchmarking and slow progress towards generalisable locomotion capabilities. Moreover, many methods are tailored to specific morphologies or tasks, offering limited transferability across embodiments and environments.
This fragmentation motivates the following central research questions: To what extent has current research achieved generalisation and robustness across diverse morphologies, terrains, and locomotion tasks? If full generalisation has not yet been realised, how can existing DRL approaches be organised and extended towards a unified framework that enables such capability in bipedal robots? In this context, the present survey seeks to categorise emerging DRL frameworks for bipedal locomotion, identify their key limitations, and outline opportunities for integration and convergence towards unification.
To address these aspects, we first clarify the ultimate goal of robot learning: to develop systems that exhibit generalisation, adaptability, and robustness across diverse morphologies, tasks, and environments. The unified framework is therefore not the final destination but a conceptual scaffold emerging from the consolidation of current DRL research efforts. Its role is to organise fragmented methodologies through shared interfaces, training conventions, and evaluation protocols, thereby promoting steady progress towards the broader goal of generalisable and adaptive robot-learning systems.
@@
Guided by these definitions, this survey examines recent advancements in DRL-based frameworks, categorising control schemes into two primary types: (i) end-to-end and (ii) hierarchical. End-to-end frameworks directly map robot states to joint-level control outputs, while hierarchical frameworks decompose decision-making into multiple layers. In hierarchical systems, a High-Level (HL) planner governs navigation and path planning, while a Low-Level (LL) controller handles fundamental locomotion tasks. The task-level decision-making tier interfaces directly with user commands or predefined tasks, forming a structured approach to robotic control.
The evolution of RL in bipedal robotics has largely advanced through the end-to-end learning paradigm. Early studies in 2004 applied simple policy-gradient methods to 2D bipeds [2004_Tedrake_RL_bipedal_policygradient, 2004_model-based_RL_learnpoincaremap_simple_bipedrobot], while later breakthroughs in DRL enabled policy training in high-fidelity physics simulators [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2018_wenhao_yu_DRL_withoutpredefine_symmetrygait_PPO_jointanglePD_]. As robotic hardware matured, an increasing variety of bipedal and humanoid platforms emerged, supporting extensive evaluation of DRL-based locomotion across diverse morphologies, as illustrated in Fig. 1. This evolution marked the transition from purely simulation-based training to simulation-to-real (sim-to-real) transfer, where policies trained in simulators are deployed on physical robots. In 2020, the first successful sim-to-real transfer of an end-to-end DRL locomotion policy was achieved on the 3D torque-controlled bipedal robot Cassie [2020_Xie_firstsim2real_]. Subsequent work explored two principal learning paradigms: reference-based learning, which leverages TO-generated data or motion-capture data to guide policy training [2021_michael_velocity-basedcontrol_motioncapturedata, 2024_speratebody_expressive_allmotion, 2024_tang_humanmimic_adversarial, 2024_Zhang_wholebody_adversarial_motion_priors]; and reference-free learning, where policies are trained entirely from scratch to autonomously discover control strategies [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie]. These developments demonstrate that end-to-end frameworks can achieve robust and versatile locomotion skills across complex terrains and dynamic environments [2022_OSUDRL_steppingstone_referencefree_predictionfeasiblefootsteps_camera_benchmark, 2023_Chongben_jumping_DDPG_multi-task_network_simulationonly, 2024_li_ucb_unifiedframework_PPO_IOput_teacherstudentcompare_IOhistory_e2etraining].
Similarly, hierarchical structures have garnered significant interest. Within this subset, the hybrid approach combines RL-based and model-based methods to enhance both planning and control strategies. Hybrid architectures often integrate learning-based and model-based modules to combine adaptability with physical consistency. One representative design couples a learned HL planner with an LL model-based controller, forming a cascade-structure or deep-planning hybrid scheme [2019_tianyu_ATRIAS_PPO_joint0level_trajectory_learnedhighlevelpolicy, 2021_duan_DRL_task-spaceaction_hiarachycotnrolscheme_inversedynamiccontroller, 2022_DRL_cascade_motionplanningpolicy_HZD_torso_PPOwithRNN]. Alternatively, DRL feedback-control hybrids embed learned control policies within model-based feedback loops to enhance tracking precision and disturbance rejection [2022_Singh_humanoid_e2erlwithfootplannar_hierarchyscheme_PPO_simulation_HRP-5P, 2023_wangsong_DRL_footsteptrakcing_hybridscheme_experiment_PPO_stair_task_LIPM_curriculumlearning]. Learned hierarchical control schemes [2017_DRL_survey] decompose locomotion into multiple layers, each focusing on specific functions such as navigation and fundamental locomotion skills [2017_Xuebinpeng_animation_deeploco_DRL_hiarachysystem_referencemotion_deepconvolutionalneuralnewtwork_jointangle_soccer., 2018_deepmimic_animation_PPO_deepmimic_desrieddirection_physics-based_animation_ATLAS_task, 2023_Wei_learned_hierarchy_framework_wheeled_bipedalrobot]. To provide a clearer overview of the current landscape, we categorise existing DRL frameworks as shown in Fig. 2.
@@
Figure 2: Classification of DRL-based control schemes. The approaches are broadly categorised into two main paradigms: end-to-end frameworks, which learn a single policy from sensory inputs to motor commands; and hierarchical frameworks, which decompose the control problem into multiple levels. Within the end-to-end paradigm, a key distinction is drawn between reference-free learning (learning from scratch) and reference-based learning (tracking a predefined motion). Hierarchical structures include hybrid control schemes, which synergistically combine learned components with traditional model-based controllers.
Current progress across both end-to-end and hierarchical paradigms indicates that a unified framework for DRL-based bipedal locomotion is still far from being realised. Establishing such a framework is essential for consolidating diverse learning pipelines, standardising evaluation metrics, and enabling transferable locomotion capabilities across different robot morphologies. As locomotion tasks become increasingly complex, ranging from basic stabilisation to dynamic parkour and loco-manipulation, the need for consistent benchmarking has intensified.The DARPA Robotics Challenge exemplified this trend by introducing one of the first large-scale evaluation platforms for bipedal humanoids performing real-world locomotion and manipulation tasks [2015_DAPRA_Challenge_1], highlighting the importance of robustness and practical deployment.
@@
Although several reviews discuss RL for general robotics [2020_systematic_review] and model-based methods for bipedal robots [2017_biped_model-based_method_review, 2021_Jenna_bipedrobot_review, 2021_Justin_review_recentporgressinleggedrobot], none specifically focus on DRL-based frameworks for bipeds. To address this gap, this survey reviews relevant literature according to the following selection criteria: (1) studies that investigate DRL frameworks specifically for bipedal robots; (2) research involving both simulated and physical bipedal robots; (3) approaches that improve policy transfer from simulation to real-world environments; and (4) publications from the last five years (2018–April 2024) sourced from reputable databases and conferences, including Google Scholar, IEEE Xplore, Web of Science, arXiv, and major robotics venues such as CoRL, RSS, ICRA, IROS, and Humanoids.
@@
The search was conducted using the key terms “deep reinforcement learning” or “reinforcement learning” in combination with “bipedal locomotion”, “bipedal walking”, “biped robot”, “humanoid robot”, or “legged robot”. The most relevant and impactful works were manually selected for further review. This survey is intended for readers with a foundational background in robotics who are transitioning to DRL methods, providing an overview of a wide range of approaches with simplified explanations where appropriate. For clarity, throughout this paper the term “humanoid robot” refers specifically to anthropomorphic bipedal robots.
@@
  • A detailed comparison of each control scheme, highlighting their strengths, limitations, and distinctive characteristics.
@@
The survey is organised as follows: Section II discusses end-to-end frameworks, categorised by learning approaches; Section III presents hierarchical frameworks, classified into three main types; Section IV outlines key limitations and challenges, linking them to the preceding discussions; Section V explores potential pathways, opportunities, and two proposed conceptual models that extend the end-to-end and hierarchical paradigms; finally, Section VI concludes the survey.
@@ II End-to-end framework
The end-to-end DRL framework represents a holistic approach in which a single neural network (NN) policy, denoted $\pi(\cdot):\mathcal{X}\rightarrow\mathcal{U}$, directly maps sensory inputs $\mathcal{X}$, such as images, LiDAR data, or proprioceptive feedback [2017_compare_actionspace_RL], together with user commands [2021_siekmann_sim2real_nonreference_perodicreward_DRL_e2e_LSTM_PPO_cassie] or pre-defined references [2021_UCB_hybridrobotics_sim2real_referencebased_HZD_gaitlibrary_e2epolicy_drl_Cassie_lowpassfilter], into joint-level control actions $\mathcal{U}$. Here, $\mathcal{X}$ represents the sensory input space, $\mathcal{U}$ refers to the space of control actions, and $\pi(\cdot)$ denotes the policy function. This framework obviates the need for manually decomposing the problem into sub-tasks, streamlining the control process.
End-to-end strategies primarily simplify the design of LL tracking to basic elements, such as a proportional–derivative (PD) controller. These methods can be broadly categorised according to their reliance on prior knowledge into two types: reference-based and reference-free. The locomotion skills developed through these diverse learning approaches exhibit considerable variation in performance and adaptability.
The following sections delve into various representation frameworks, exploring their characteristics, limitations, and strengths in comprehensive detail. To facilitate an understanding of these distinctions, Table I provides a succinct overview of the frameworks discussed.
TABLE I: Summary and comparison of reference-based and reference-free learning approaches for the end-to-end framework. The dashed line in the implementation flow chart indicates optional steps.00footnotetext: Forward Walk involves bipeds walking straight ahead. Unidirectional Walk enables bipeds to move forward and backward within a range of desired velocities. Omnidirectional Walk grants bipeds the ability to walk in any direction. Versatile Walk allows the bipeds to walk forward, backward, turn, and move sideways, providing extensive movement capabilities. Periodic Motions entails the execution of various repeated gait patterns, such as walking, hopping, or galloping. Versatile Jump refers to jumping towards different desired targets. Versatile Motions cover performing a broad array of motions, both periodic and aperiodic such as jumping.